A 150N force acts on a 300kg car calculate how long it takes this force to change the car velocity from 2m/s to 6m/s
1
Expert's answer
2020-01-27T10:00:58-0500

We can find the time taken to change the car velocity from 2m/s to 6m/s from the definition of the impulse:


"F \Delta t = \Delta p = m \Delta v,""F \Delta t = m(v_f - v_i),"

here, "F" is the force acting on the car, "\Delta t" is the amount of time in which the force is acting (or the time taken to change the car velocity), "\Delta p" is the impulse change, "m" is the mass of the car, "\Delta v" is the rate of change of car's velocity, "v_i" is the initial velocity of the car and "v_f" is the final velocity of the car.

From this equation we can find the time taken to change the car velocity:


"\Delta t = \dfrac{m(v_f - v_i)}{F} = \dfrac{300kg \cdot (6 \dfrac{m}{s} - 2 \dfrac{m}{s})}{150N} = 8 s."

Answer:

"\Delta t = 8 s."